# O. Al-Ghattas, J. Chen, D. Sanz-Alonso

### v1

@@ Abstract
This paper establishes sharp dimension-free concentration inequalities and expectation bounds for the deviation of the sum of simple random tensors from its expectation. As part of our analysis, we use generic chaining techniques to obtain a sharp high-probability upper bound on the suprema of multi-product empirical processes. In so doing, we generalize classical results for quadratic and product empirical processes to higher-order settings.
@@ 1 Introduction
In addition, Theorem 2.1 establishes sharp high-probability and expectation bounds for sub-Gaussian data. To prove Theorem 2.1, we analyze the suprema of a multi-product empirical process. In our second main result, Theorem 2.2, we leverage generic chaining techniques to obtain a sharp high-probability bound on
@@ 1.1 Outline
Section 2 states our two main results and discusses related work. The study of concentration of simple random tensors and multi-product empirical processes is carried out in Sections 3 and 4, respectively. Section 5 closes with conclusions and future directions.
@@ 2 Main results
Here we state the two main theorems of the paper and compare them with existing results. Subsection 2.1 focuses on the concentration of simple random tensors and Subsection 2.2 focuses on multi-product empirical processes.
@@ 2.1 Concentration of simple random tensors
[…]
We will focus on Gaussian and sub-Gaussian random variables. Recall that an $H$-valued random variable $X$ is called (sub)-Gaussian iff all one-dimensional projections $\langle X,v\rangle$ for $v\in H$ are (sub)-Gaussian real-valued random variables. Furthermore, $X$ is called pre-Gaussian iff there exists a centered Gaussian random variable $Y$ in $H$ with the same covariance operator as that of $X.$
@@
The remainder of this subsection provides detailed comparisons between Theorem 2.1 and existing results. In Remark 2.1, we compare with concentration results for sums of simple tensors, while in Remark 2.2 we discuss the special case $p=2.$
@@
$\bigg{\|}\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}-\mathbb{E}\,X^{\otimes p}\bigg{\|}=\sup_{\|v\|=1}\bigg{|}\frac{1}{N}\sum_{i=1}^{N}\left\langle X_{i},v\right\rangle^{p}-\mathbb{E}\langle X,v\rangle^{p}\bigg{|}\overset{\text{($\star$)}}{\lesssim_{p}}\|\Sigma\|^{p/2}\sqrt{\frac{r(\Sigma)}{N}}.$
@@
2.2 Multi-product empirical processes
This subsection contains the statement of our second main result, Theorem 2.2, which gives a general upper bound on the suprema of multi-product empirical processes. Theorem 2.2 will be key to establish our upper bounds in Theorem 2.1. Let $X,X_{1},\ldots,X_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$ be a sequence of random variables on a probability space $(\Omega,\mu)$. For $p\geq 1$ (not necessarily an integer), we consider the multi-product centered empirical process indexed by a class $\mathcal{F}$ of functions on $(\Omega,\mu)$ given by
@@ 2.2 Multi-product empirical processes
Assume $0\in\mathcal{F}$ or $\mathcal{F}$ is symmetric (i.e., $f\in\mathcal{F}$ implies $-f\in\mathcal{F}$). For any $p\geq 2$ and $u\geq 1$, it holds with probability at least $1-\exp(-u^{2}(\gamma(\mathcal{F},\psi_{2})/d_{\psi_{2}}(\mathcal{F}))^{2})$ that
@@ 3 Concentration of simple random tensors
where :=$\mathcal{F}:=\left\{\langle\cdot,v\rangle:v\in U_{H}\right\},U_{H}:=\left\{v\in H:\|v\|=1\right\}$. It is clear that $\mathcal{F}$ is symmetric. Since $X$ is sub-Gaussian, the $\psi_{2}$-norm of linear functionals is equivalent to the $L_{2}$-norm. Hence,
@@
[…]
@@
()${\geq}\frac{c^{\prime}(p)}{\sqrt{N}}\bigg{(}\|\Sigma\|^{(p-1)/2}\mathbb{E}\|X\|-\sqrt{\frac{2}{\pi}}\|\Sigma\|^{p/2}\bigg{)}$
[…]
@@
4 Multi-product empirical processes
This section studies multi-product empirical processes
@@ 4.1 Background and auxiliary results
According to Lemma 4.2, the influence of $z$ on the Bernoulli process depends on the $\ell_{1}$-norm of the largest $k$ coordinates of $(|z_{i}|)_{i=1}^{N}$ and the $\ell_{2}$-norm of the smallest $N-k$ coordinates. As noted in [51], the graded $L_{q}$ norm plays a key role in analyzing the monotone non-increasing rearrangement of $N$ independent copies of $Z$. In particular, [51, Corollary 3.6] provides a useful way to decompose a collection of random vectors into their largest and smallest coordinate components, a technique central to [51]. Below, we present the statement obtained by setting $q=2r$ and $\beta=1/2$ in their notation.
@@
Let $(\mathcal{F}_{s})_{s\geq 0}\subset\mathcal{F}$ be an admissible sequence. For $f\in\mathcal{F}$ and $s\geq 0$, we define $\Delta_{s}f:=\pi_{s+1}f-\pi_{s}f$. To analyze multi-product empirical processes of order $p$, we apply Lemma 4.3 to the classes $\mathcal{F}_{s}$ and $\{\Delta_{s}f:f\in\mathcal{F}\}$ with carefully chosen values of $r$ that depend on $p$, leading to the following corollary.
@@
By definition of an admissible sequence, we have $|\mathcal{F}_{s}|\leq 2^{2^{s}}.$ Further, $|\mathcal{F}_{s}|\cdot|\mathcal{F}_{s+1}|\leq 2^{2^{s}}\cdot 2^{2^{s+1}}\leq 2^{2^{s+2}},$ which implies that for $\Delta_{s}f=\pi_{s+1}f-\pi_{s}f$, |$|\{\Delta_{s}f:f\in\mathcal{F}\}|\leq 2^{2^{s+2}}$. Therefore, the result follows by applying Lemma 4.3, for every $s\geq s_{0}$, to the class $\mathcal{F}_{s}$ with the choice $r=4(p-1)\geq 1$ and to the class $\{\Delta_{s}f:f\in\mathcal{F}\}$ with $r=4$, and taking a union bound to get the probability estimate
@@
We demonstrate the optimality of the bound (4.1) via a simple example. Let $H=\mathbb{R}^{d}$ and consider a symmetric compact set $K\subset\mathbb{R}^{d}$. Define the function class =$\mathcal{F}=\left\{\langle\cdot,v\rangle:v\in K\right\}$, which consists of linear functions indexed by $K$. Let $X_{1},\ldots,X_{N}$ be i.i.d. standard Gaussian random vectors in $\mathbb{R}^{d}$, and let $v^{*}\in K$ be a point attaining the maximal Euclidean norm on $\mathbb{R}^{d}$, i.e., $\|v^{*}\|_{\ell_{2}}=\sup_{v\in K}\|v\|_{\ell_{2}}$. Observe that
$\mathbb{E}\sup_{f\in\mathcal{F}}\bigg{(}\sum_{i=1}^{N}|f|^{m}\left(X_{i}\right)\bigg{)}^{1/m}=\mathbb{E}\sup_{v\in K}\bigg{(}\sum_{i=1}^{N}|\langle X_{i},v\rangle|^{m}\bigg{)}^{1/m}$ ${\gtrsim_{m}}\mathbb{E}\sup_{v\in K}|\langle X_{1},v\rangle|+N^{1/m}\|v^{*}\|_{\ell_{2}}$
[…]
@@
We first observe that the condition $0\in\mathcal{F}$ or the symmetry of $\mathcal{F}$ (i.e., $f\in\mathcal{F}$ implies $-f\in\mathcal{F}$) ensures that $\gamma(\mathcal{F},\psi_{2})\gtrsim d_{\psi_{2}}(\mathcal{F})$ (Lemma 4.6), a fact that we will use repeatedly in the proof.
@@
()${\leq}p\sum_{i\in I_{s}}|(\Delta_{s}f)_{i}|\left(|\pi_{s+1}f|^{p-1}+|\pi_{s}f|^{p-1}\right)_{i}$
@@
Observe that |$\left|\left\{\pi_{s}f:f\in\mathcal{F}\right\}\right|\leq|\mathcal{F}_{s}|\leq 2^{2^{s}}$, |$\left|\left\{\pi_{s+1}f:f\in\mathcal{F}\right\}\right|\leq|\mathcal{F}_{s+1}|\leq 2^{2^{s+1}}$ and |$\left|\left\{\Delta_{s}f:f\in\mathcal{F}\right\}\right|\leq|\mathcal{F}_{s}|\cdot|\mathcal{F}_{s+1}|\leq 2^{2^{s}}\cdot 2^{2^{s+1}}\leq 2^{2^{s+2}}$. For $u\geq 4$, we sum the probabilities for all $s\geq s_{0}$ to obtain that, with $(\varepsilon_{i})_{i=1}^{N}$ probability at least
@@
()${\lesssim}_{p}\sup_{f\in\mathcal{F}}\bigg{(}\sum_{i=1}^{N}f^{2(p-1)}_{i}\bigg{)}^{1/2}u\widetilde{\Lambda}_{s_{0},cu}(\mathcal{F})+uN^{1/2}d^{p-1}_{\psi_{2}}(\mathcal{F})\bigg{(}\sum_{s\geq s_{0}}2^{s/2}\|\Delta_{s}f\|_{\psi_{2}}\bigg{)}+u2^{s_{0}/2}N^{1/2}d^{p}_{\psi_{2}}(\mathcal{F})$
[…]
@@
where ()(*)( ∗ ) is some random variable, $C_{1}(p)\geq 1,C_{2}(p)\geq 1,c_{3}(p)>0,\omega\geq 1,A\geq 0,B\geq 0$ are constants. Then, for any $u\geq 1$, it holds with probability at least $1-\exp(-u^{2}\omega)$ that
@@
This subsection presents the proof of Theorem 4.5, which establishes a uniform bound on the $\ell_{m}$-norm of the coordinate projection vector of the function class $\mathcal{F}$. The proof leverages generic chaining and builds upon ideas from Bednorz [11, Section 5] for analyzing quadratic processes, while also addressing new challenges that arise when handling arbitrary $m\geq 1$. A key component of our approach is the use of a different Young function, defined as $\varphi(x)=2^{\min\{(\sqrt{N}x)^{2/m},\,x^{2}\}}-1$, along with $\alpha$-sub-exponential inequalities [22] to precisely characterize the tail behavior. This leads to a distinct cut-off point in the chaining process —depending on $m$— that separates sub-Gaussian and sub-exponential regimes. Similar cut-off phenomena between these two regimes first appeared in the analysis of quadratic empirical processes (see, e.g., [48, Section 5], [17, Theorem 5.5] and [70, Theorem 14.2.6]). However, our primary focus here is to control the non-centralized $\ell_{m}$-norm of the coordinate projection vector for any $m\geq 1$.
@@
()${\lesssim}_{m}\varphi^{-1}(N_{s_{k+1}})+\varphi^{-1}\bigg{(}\frac{1}{N_{s_{k+1}}^{3}}\sum_{g,h\in\mathcal{F}_{s_{k+1}}}\varphi\left(|\mathsf{X_{1}}(g,h)|\right)\bigg{)}$
@@
()${\lesssim}_{m}N^{-1/m}\gamma(\mathcal{F},\psi_{2})+\|f-\pi_{s^{*}}f\|_{\psi_{2}}\left(1+N^{-1/2m}(\varphi^{-1}(Z_{1}))^{1/m}\right)$
@@
()${\leq}\sum_{k=0}^{\infty}2^{s_{k+1}/2}\left(\|\pi_{s_{k+1}}f-f\|_{\psi_{2}}+2\|\pi_{s_{k+1}-1}f-f\|_{\psi_{2}}\right)$

### v2

@@ Abstract
This paper establishes sharp dimension-free concentration inequalities and expectation bounds for the deviation of the sum of simple random tensors from its expectation. As part of our analysis, we use generic chaining techniques to obtain a sharp high-probability upper bound on the suprema of $L_{p}$ empirical processes. In so doing, we generalize classical results for quadratic and product empirical processes to higher-order settings.
@@ 1 Introduction
In addition, Theorem 2.1 establishes sharp high-probability and expectation bounds for sub-Gaussian data. To prove Theorem 2.1, we analyze the suprema of a $L_{p}$ empirical process. In our second main result, Theorem 2.2, we leverage generic chaining techniques to obtain a sharp high-probability bound on
@@ 1.1 Outline
Section 2 states our two main results and discusses related work. The study of concentration of simple random tensors and $L_{p}$ empirical processes is carried out in Sections 3 and 4, respectively. Section 5 closes with conclusions and future directions.
@@ 2 Main results
Here we state the two main theorems of the paper and compare them with existing results. Subsection 2.1 focuses on the concentration of simple random tensors and Subsection 2.2 focuses on $L_{p}$ empirical processes.
@@ 2.1 Concentration of simple random tensors
where $\|\cdot\|$ denotes the norm in the tensor space $H^{\otimes p}.$ Notice that for $p=2,$ the operator norm of the covariance operator $\Sigma:H\to H$ agrees with the norm of the covariance $\mathbb{E}\,X^{\otimes 2}$ as a bilinear form in $H^{\otimes 2}.$ Consequently, we will slightly abuse notation and denote by $\|\Sigma\|$ the operator norm of the covariance operator. Furthermore, to simplify the notation, we will henceforth omit the subscript $H$ in the inner-product $\langle\cdot,\cdot\rangle_{H}$ and norm $\|\cdot\|_{H}$.
We will focus on Gaussian and sub-Gaussian random variables. Recall that an $H$-valued random variable $X$ is called Gaussian iff all one-dimensional projections $\langle X,v\rangle$ for $v\in H$ are Gaussian real-valued random variables. A centered random variable $X$ in $H$ is called sub-Gaussian iff
$\|\langle X,v\rangle\|_{\psi_{2}}\lesssim\|\langle X,v\rangle\|_{L_{2}},\qquad\forall\ v\in H,$
where the Orlicz $\psi_{2}$-norm of a real-valued random variable $Y$ is defined as $\|Y\|_{\psi_{2}}:=\inf\left\{c>0:\mathbb{E}\exp(Y^{2}/c^{2})\leq 2\right\}$. Furthermore, $X$ is called pre-Gaussian iff there exists a centered Gaussian random variable $Y$ in $H$ with the same covariance operator as that of $X.$
@@
For sub-Gaussian and pre-Gaussian random variables satisfying that, for fixed $K>0$,
$\|\langle X,v\rangle\|_{\psi_{2}}\leq K\|\langle X,v\rangle\|_{L_{2}},\qquad\forall\ v\in H,$ (2.2)
a straightforward modification of the proof of Theorem 2.1 yields
$\mathbb{E}\bigg\|\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}-\mathbb{E}\,X^{\otimes p}\bigg\|\lesssim_{p}K^{p}\|\Sigma\|^{p/2}\bigg(\sqrt{\frac{r(\Sigma)}{N}}+\frac{r(\Sigma)^{p/2}}{N}\bigg),$
as well as a similar high-probability upper bound. For a centered Gaussian random variable, (2.2) holds with $K=\sqrt{8/3}$. When $p=2$, the dependence on $K^{2}$ is consistent with the literature on covariance estimation; see [78, Theorem 4.7.1]. For $p>2$, the dependence on $K^{p}$ agrees with the result of [84]. The notation $\lesssim_{p}$ indicates that the implicit constant depends on the tensor order $p$, and this dependence is generally super-exponential in $p$. This can be seen from the fact that, when $p$ is even and $X\sim\mathcal{N}(0,\Sigma)$,
$\|\mathbb{E}\,X^{\otimes p}\|=(p-1)!!\|\Sigma\|^{p/2},\quad\text{ and }\quad(p-1)!!\asymp\sqrt{2(p-1)}\Big(\frac{p-1}{e}\Big)^{\frac{p-1}{2}}\asymp e^{\frac{1}{2}p\log p}.$
In Theorem 2.1 and in the comparisons that we will provide in the remainder of this subsection, we assume $K\asymp 1$ for simplicity. In Remark 2.1, we compare with concentration results for sums of simple tensors, while in Remark 2.2 we discuss the special case $p=2.$
@@
$\bigg\|\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}-\mathbb{E}\,X^{\otimes p}\bigg\|=\sup_{\|v\|=1}\bigg|\frac{1}{N}\sum_{i=1}^{N}\left\langle X_{i},v\right\rangle^{p}-\mathbb{E}\langle X,v\rangle^{p}\bigg|\overset{\hskip-2.84544pt\text{($\star$)}}{\lesssim_{p}}\|\Sigma\|^{p/2}\sqrt{\frac{r(\Sigma)}{N}}.$
@@
2.2 $L_{p}$ empirical processes
This subsection contains the statement of our second main result, Theorem 2.2, which gives a general upper bound on the suprema of $L_{p}$ empirical processes. Theorem 2.2 will be key to establish our upper bounds in Theorem 2.1. Let $X,X_{1},\ldots,X_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$ be a sequence of random variables on a probability space $(\Omega,\mu)$. For $p\geq 1$ (not necessarily an integer), we consider the centered $L_{p}$ empirical process indexed by a class $\mathcal{F}$ of functions on $(\Omega,\mu)$ given by
@@ 2.2 $L_{p}$ empirical processes
Assume $0\in\mathcal{F}$ or $\mathcal{F}$ is symmetric (i.e., $f\in\mathcal{F}$ implies $f\in\mathcal{F}$). For any $p\geq 2$ and $u\geq 1$, it holds with probability at least $1-\exp(-u^{2}(\gamma(\mathcal{F},\psi_{2})/d_{\psi_{2}}(\mathcal{F}))^{2})$ that
@@ 3 Concentration of simple random tensors
where $\mathcal{F}:=\left\{\langle\cdot,v\rangle:v\in U_{H}\right\},U_{H}:=\left\{v\in H:\|v\|=1\right\}$. It is clear that $\mathcal{F}$ is symmetric. Since $X$ is sub-Gaussian, the $\psi_{2}$-norm of linear functionals is equivalent to the $L_{2}$-norm. Hence,
@@
$\overset{\text{($\star$)}}{\geq}\frac{1}{N}\left(\mathbb{E}\|X_{1}\|^{p}-C(p)\|\Sigma\|^{p/2}\right)\overset{\text{($\star\star$)}}{\geq}\frac{1}{N}\left(\mathrm{Tr}(\Sigma)^{p/2}-C(p)\|\Sigma\|^{p/2}\right)=\frac{1}{N}\|\Sigma\|^{p/2}\left(r(\Sigma)^{p/2}-C(p)\right).$ (3.3)
Here, ($\star$) follows by $\mathbb{E}_{X_{1}}\left[\sup_{\|v\|=1}|\langle X_{1},v\rangle^{p}|\right]=\mathbb{E}\|X_{1}\|^{p}$ and $\sup_{\|v\|=1}|\mathbb{E}\langle X,v\rangle^{p}|\leq C(p)\|\Sigma\|^{p/2}$, since $\langle X,v\rangle\overset{\mathsf{d}}{\sim}\mathcal{N}(0,\langle\Sigma v,v\rangle)$, while ($\star\star$) follows by Jensen’s inequality $\mathbb{E}\|X_{1}\|^{p}\geq\left(\mathbb{E}\|X_{1}\|^{2}\right)^{p/2}=\mathrm{Tr}(\Sigma)^{p/2}$ for $p\geq 2$.
@@
$\overset{\text{($\star$)}}{\geq}\frac{c^{\prime}(p)}{\sqrt{N}}\bigg(\|\Sigma\|^{(p-1)/2}\mathbb{E}\|X\|-\sqrt{\frac{2}{\pi}}\|\Sigma\|^{p/2}\bigg)$
$\overset{\text{($\star\star$)}}{\geq}c^{\prime}(p)\|\Sigma\|^{p/2}\bigg(\frac{c^{\prime}\sqrt{r(\Sigma)}-\sqrt{2/\pi}}{\sqrt{N}}\bigg),$ (3.9)
where ($\star$) follows by taking the eigenvector of $\Sigma$ corresponding to the largest eigenvalue, and ($\star\star$) follows since for a Gaussian vector $X$, $\mathbb{E}\|X\|\asymp\sqrt{\mathbb{E}\|X\|^{2}}=\sqrt{\mathrm{Tr}(\Sigma)}$.
@@
4 $L_{p}$ empirical processes
This section studies $L_{p}$ empirical processes
@@ 4.1 Background and auxiliary results
According to Lemma 4.2, the influence of $z$ on the Bernoulli process depends on the $\ell_{1}$-norm of the largest $k$ coordinates of $(|z_{i}|)_{i=1}^{N}$ and the $\ell_{2}$-norm of the smallest NkN-k coordinates. As noted in [51], the graded $L_{q}$ norm plays a key role in analyzing the monotone non-increasing rearrangement of $N$ independent copies of $Z$. In particular, [51, Corollary 3.6] provides a useful way to decompose a collection of random vectors into their largest and smallest coordinate components, a technique central to [51]. Below, we present the statement obtained by setting $q=2r$ and $\beta=1/2$ in their notation.
@@
Let $(\mathcal{F}_{s})_{s\geq 0}\subset\mathcal{F}$ be an admissible sequence. For $f\in\mathcal{F}$ and $s\geq 0$, we define $\Delta_{s}f:=\pi_{s+1}f-\pi_{s}f$. To analyze $L_{p}$ empirical processes, we apply Lemma 4.3 to the classes $\mathcal{F}_{s}$ and $\{\Delta_{s}f:f\in\mathcal{F}\}$ with carefully chosen values of $r$ that depend on $p$, leading to the following corollary.
@@
By definition of an admissible sequence, we have $|\mathcal{F}_{s}|\leq 2^{2^{s}}.$ Further, $|\mathcal{F}_{s}|\cdot|\mathcal{F}_{s+1}|\leq 2^{2^{s}}\cdot 2^{2^{s+1}}\leq 2^{2^{s+2}},$ which implies that for $\Delta_{s}f=\pi_{s+1}f-\pi_{s}f$, $|\{\Delta_{s}f:f\in\mathcal{F}\}|\leq 2^{2^{s+2}}$. Therefore, the result follows by applying Lemma 4.3, for every $s\geq s_{0}$, to the class $\mathcal{F}_{s}$ with the choice $r=4(p-1)\geq 1$ and to the class $\{\Delta_{s}f:f\in\mathcal{F}\}$ with $r=4$, and taking a union bound to get the probability estimate
@@
We demonstrate the optimality of the bound (4.1) via a simple example. Let $H=\mathbb{R}^{d}$ and consider a symmetric compact set $K\subset\mathbb{R}^{d}$. Define the function class $\mathcal{F}=\left\{\langle\cdot,v\rangle:v\in K\right\}$, which consists of linear functions indexed by $K$. Let $X_{1},\ldots,X_{N}$ be i.i.d. standard Gaussian random vectors in $\mathbb{R}^{d}$, and let $v^{*}\in K$ be a point attaining the maximal Euclidean norm on $\mathbb{R}^{d}$, i.e., $\|v^{*}\|_{\ell_{2}}=\sup_{v\in K}\|v\|_{\ell_{2}}$. Observe that
$\mathbb{E}\sup_{f\in\mathcal{F}}\bigg(\sum_{i=1}^{N}|f|^{m}\left(X_{i}\right)\bigg)^{1/m}=\mathbb{E}\sup_{v\in K}\bigg(\sum_{i=1}^{N}|\langle X_{i},v\rangle|^{m}\bigg)^{1/m}$ $\overset{\hskip-2.84544pt\text{($\star$)}}{\gtrsim_{m}}\mathbb{E}\sup_{v\in K}|\langle X_{1},v\rangle|+N^{1/m}\|v^{*}\|_{\ell_{2}}$
$\overset{\text{($\star\star$)}}{\asymp}\gamma(\mathcal{F},\psi_{2})+N^{1/m}d_{\psi_{2}}(\mathcal{F}),$
where ($\star$) follows by separately analyzing the case $N=1$ and the contribution of the maximal element $v^{*}\in K$, and ($\star\star$) follows by Talagrand’s majorizing-measure theorem [70, Theorem 2.10.1], which gives $\mathbb{E}\sup_{v\in K}\langle X_{1},v\rangle\asymp\gamma(\mathcal{F},\psi_{2})$, along with the relation $\|v^{*}\|_{\ell_{2}}=\sup_{v\in K}\|v\|_{\ell_{2}}\asymp\sup_{f\in\mathcal{F}}\|f\|_{\psi_{2}}=d_{\psi_{2}}(\mathcal{F})$. This confirms the necessity of the deterministic terms in (4.1). Furthermore, note that
@@
We first observe that the condition $0\in\mathcal{F}$ or the symmetry of $\mathcal{F}$ (i.e., $f\in\mathcal{F}$ implies $f\in\mathcal{F}$) ensures that $\gamma(\mathcal{F},\psi_{2})\gtrsim d_{\psi_{2}}(\mathcal{F})$ (Lemma 4.6), a fact that we will use repeatedly in the proof.
@@
$\overset{\text{($\star$)}}{\leq}p\sum_{i\in I_{s}}|(\Delta_{s}f)_{i}|\left(|\pi_{s+1}f|^{p-1}+|\pi_{s}f|^{p-1}\right)_{i}$
@@
Observe that $\left|\left\{\pi_{s}f:f\in\mathcal{F}\right\}\right|\leq|\mathcal{F}_{s}|\leq 2^{2^{s}}$, $\left|\left\{\pi_{s+1}f:f\in\mathcal{F}\right\}\right|\leq|\mathcal{F}_{s+1}|\leq 2^{2^{s+1}}$ and $\left|\left\{\Delta_{s}f:f\in\mathcal{F}\right\}\right|\leq|\mathcal{F}_{s}|\cdot|\mathcal{F}_{s+1}|\leq 2^{2^{s}}\cdot 2^{2^{s+1}}\leq 2^{2^{s+2}}$. For $u\geq 4$, we sum the probabilities for all $s\geq s_{0}$ to obtain that, with $(\varepsilon_{i})_{i=1}^{N}$ probability at least
@@
$\overset{\text{($\star$)}}{\lesssim}_{p}\sup_{f\in\mathcal{F}}\bigg(\sum_{i=1}^{N}f^{2(p-1)}_{i}\bigg)^{1/2}u\widetilde{\Lambda}_{s_{0},cu}(\mathcal{F})+uN^{1/2}d^{p-1}_{\psi_{2}}(\mathcal{F})\bigg(\sum_{s\geq s_{0}}2^{s/2}\|\Delta_{s}f\|_{\psi_{2}}\bigg)+u2^{s_{0}/2}N^{1/2}d^{p}_{\psi_{2}}(\mathcal{F})$
$\overset{\text{($\star\star$)}}{\lesssim}_{p}\sup_{f\in\mathcal{F}}\bigg(\sum_{i=1}^{N}f^{2(p-1)}_{i}\bigg)^{1/2}u\widetilde{\Lambda}_{s_{0},cu}(\mathcal{F})+uN^{1/2}d^{p-1}_{\psi_{2}}(\mathcal{F})\left(\gamma(\mathcal{F},\psi_{2})+2^{s_{0}/2}d_{\psi_{2}}(\mathcal{F})\right),$
where ($\star$) follows by choosing an almost optimal admissible sequence in $\mathcal{F}$ and using the definition of $\widetilde{\Lambda}_{s_{0},u}(\mathcal{F})$, and the fact that $\|\pi_{s}f\|_{L_{8(p-1)}}^{p}\lesssim_{p}\|\pi_{s}f\|_{\psi_{2}}^{p}\leq d_{\psi_{2}}^{p}(\mathcal{F}),\|\pi_{s_{0}}f\|_{L_{8(p-1)}}^{p}\lesssim_{p}\|\pi_{s_{0}}f\|_{\psi_{2}}^{p}\leq d_{\psi_{2}}^{p}(\mathcal{F})$; ($\star\star$) follows by
@@
where ()(*) is some random variable, $C_{1}(p)\geq 1,C_{2}(p)\geq 1,c_{3}(p)>0,\omega\geq 1,A\geq 0,B\geq 0$ are constants. Then, for any $u\geq 1$, it holds with probability at least $1-\exp(-u^{2}\omega)$ that
@@
This subsection presents the proof of Theorem 4.5, which establishes a uniform bound on the $\ell_{m}$-norm of the coordinate projection vector of the function class $\mathcal{F}$. The proof leverages generic chaining and builds upon ideas from Bednorz [11, Section 5] for analyzing quadratic processes, while also addressing new challenges that arise when handling arbitrary $m\geq 1$. A key component of our approach is the use of a different Young function, defined as $\varphi(x)=2^{\min\{(\sqrt{N}x)^{2/m},\,x^{2}\}}-1$, along with $\alpha$-sub-exponential inequalities [22] to precisely characterize the tail behavior. This leads to a distinct cut-off point in the chaining process —depending on mm— that separates sub-Gaussian and sub-exponential regimes. Similar cut-off phenomena between these two regimes first appeared in the analysis of quadratic empirical processes (see, e.g., [48, Section 5], [17, Theorem 5.5] and [70, Theorem 14.2.6]). However, our primary focus here is to control the non-centralized $\ell_{m}$-norm of the coordinate projection vector for any $m\geq 1$.
@@
$\overset{\text{($\star$)}}{\lesssim}_{m}\varphi^{-1}(N_{s_{k+1}})+\varphi^{-1}\bigg(\frac{1}{N_{s_{k+1}}^{3}}\sum_{g,h\in\mathcal{F}_{s_{k+1}}}\varphi\left(|\mathsf{X_{1}}(g,h)|\right)\bigg)$
@@
$\overset{\text{($\star$)}}{\lesssim}_{m}N^{-1/m}\gamma(\mathcal{F},\psi_{2})+\|f-\pi_{s^{*}}f\|_{\psi_{2}}\left(1+N^{-1/2m}(\varphi^{-1}(Z_{1}))^{1/m}\right)$
@@
$\overset{\text{($\star$)}}{\leq}\sum_{k=0}^{\infty}2^{s_{k+1}/2}\left(\|\pi_{s_{k+1}}f-f\|_{\psi_{2}}+2\|\pi_{s_{k+1}-1}f-f\|_{\psi_{2}}\right)$